### v1

# Information Scrambling in ̵Bosonic Gaussian Dynamics

Abstract
We show that randomness in quadratic Hamiltonians for multimode bosonic systems results in certain information scrambling diagnostics, mirroring those seen in chaotic systems. Specifically, considering initial Gaussian states, we observe the disappearance of the memory effect in entanglement dynamics and the negative value of tripartite mutual information. However, in contrast to chaotic systems, we find that the spectral form factor exhibits a non-linear ramp and the out-of-time-ordered correlators display a power law growth. These results show that information scrambling, associated with randomness, is a distinct feature from quantum chaos. Moreover, our results provide insight into the dynamics of Gaussian states of continuous-variable systems that are useful and available resources for quantum information processing.
Introduction.
Originating from the black-hole information paradox, scrambling of quantum information has become a central feature differentiating between integrable systems and chaotic systems [1, 2, 3]. Several diagnostics have been proposed to identify information scrambling such as the spectral form factor (SSF) [4, 5, 6], operator entanglement [7, 8, 9, 10, 11], entanglement spread [12, 13, 14, 15, 16, 17, 18, 19, 20], out-of-time-ordered correlators (OTOC) [21, 3, 22, 23, 24] and tripartite mutual information (TMI) [25]. However, it has recently been shown that observing an information scrambling diagnostic does not necessarily imply quantum chaos in the system. Specifically, non-chaotic/integrable systems have been identified that exhibit the exponential growth of OTOC [26, 27, 28] and operator entanglement [29].
These recent observations raise the question of how quantum information scrambling, distinct from quantum chaos, should be characterized in quantum systems. In particular, identifying the features that drive information scrambling can provide insight into the dynamics of complex quantum systems.
In this paper, we investigate quantum information scrambling in multimode bosonic systems with quadratic Hamiltonians. These systems are integrable however, with a sufficient amount of randomness in the Hamiltonian, admit single-particle chaos [30, 31], i.e. the single-particle sector of the Hilbert space is characterized by random-matrix-like level statistics. We examine different amounts of randomness, ranging from the local translational invariant to the completely random non-local models  [^1] The latter model may be seen as bosonic counterparts for the SYK2 model.. For initial Gaussian states and with even a small amount of randomness, preserving the locality but removing translational invariance in the dynamics, we observe the vanishing of the memory effect during entanglement spread and the negative values of TMI. These results imply scrambling of information in these systems, similar to what has been predicted for holographic states [17, 25].
We also consider the SFF and OTOC as other information scrambling diagnostics for these systems. We observe that the SFF exhibits a dip followed by a ramp, attributed to the single-particle sector of the Hamiltonian. However, we find non-linear ramps, deviating from the expected behavior in chaotic systems. Additionally, we show that the OTOC displays a power law growth, distinct from the exponential growth seen in chaotic systems. Consequently, these diagnostic tools remain unaffected by the randomness added to the dynamics of the system.
Our study not only sheds light on new aspects of information scrambling but also offers new insights into the dynamics of continuous-variable systems. Gaussian states of these systems are particularly intriguing because of their experimental accessibility. Specifically, quadratic Hamiltonians for bosonic systems lead to Gaussian dynamics, which can be efficiently simulated using linear optical circuits [33]. These circuits are readily accessible and have been used in Gaussian boson sampling experiments [34, 35, 36, 37]. Therefore, our results can be experimentally verified and applied to investigate the presence of randomness in the dynamics of such systems.
This paper is structured as follows. We first explain the setup, including Gaussian states and quadratic Hamiltonian models that we use. Then, we report our results on the dynamics of entanglement, TMI, SFF, and OTOC in the presence of randomness in the Hamiltonian.
Setup.
We consider $N$-mode bosonic systems described by the vector of quadrature operators $r=(q_{1},\dots,q_{N},p_{1},\dots,p_{N})^{\top}$, satisfying the canonical commutation relations $[q_{j},p_{k}]=i\delta_{jk}$ ($\hbar=1$). Quantum states are described by density operators $\rho$ ($\rho\succeq 0$ and ${\rm tr}(\rho)=1$), and those that can be represented by a Gaussian Wigner function can represent are known as Gaussian states [33, 38]. These states can be uniquely characterized using the first-order moment vector $\langle r\rangle={\rm tr}(\rho r)$ and the covariance matrix with the matrix elements given by $\bm{\sigma}_{jk}=\langle r_{j}r_{k}+r_{k}r_{j}\rangle-2\langle r_{j}\rangle\langle r_{k}\rangle$. The covariance matrix is symmetric and positive and satisfies the uncertainty relation $\bm{\sigma}+i\bm{J}\geq 0$, where $\bm{J}$ is the symplectic form whose matrix elements are given by $\bm{J}_{jk}=i[r_{k},r_{j}]$. Reduced states of a Gaussian state with $\bm{\sigma}$ are Gaussian whose covariance matrices are submatrices of $\bm{\sigma}$. We consider states with $\langle r\rangle=0$, as this condition can always be satisfied by using local unitary displacement operations. A special example is the product thermal state of $N$-mode system whose $\langle r\rangle=0$ and $\bm{\sigma}_{\rm{th}}=\text{diag}(\nu_{1},\dots,\nu_{N},\nu_{1},\dots,\nu_{N})$, where $\nu_{k}\geq 1$ with $\nu_{k}=1$ corresponding to the vacuum state in the $k$th mode.
Unitary transformations that preserve the Gaussianity of quantum states are known as Gaussian unitaries. Up to local displacement operations, Gaussian unitaries are associated with quadratic Hamiltonians of this form
$H=\frac{1}{2}r^{\top}\bm{M}r$ (1)
with matrix $\bm{M}$ being real, symmetric, and positive definite  [^2] In general, $\bm{M}$ can be a Hermitian matrix, which can then be written as a sum of real symmetric and anti-symmetric matrices, $\bm{M}=\bm{M}_{\rm{sym}}+i\bm{M}_{\rm{asy}}$. However, the anti-symmetric part $\bm{M}_{\rm{asy}}$ only contributes an overall phase that can be ignored.. Such Gaussian unitary can be described by $e^{iHt}re^{-iHt}=\bm{S}_{t}r$, where $\bm{S}=e^{\bm{J}\bm{M}t}$ is a symplectic matrix, satisfying $\bm{S}\bm{J}\bm{S}^{\top}=\bm{J}$. Using this, one can see that the evolution of Gaussian states under Gaussian unitaries, $\rho_{t}=e^{-iHt}\rho e^{iHt}$, can be described in terms of symplectic transformations on the covariance matrices, $\bm{\sigma}_{t}=\bm{S}\bm{\sigma}\bm{S}^{\top}$. Gaussian unitaries with orthogonal symplectic matrices, preserve the mean energy of the system and are known as passive transformations. In quantum optics, these transformations can be realized using lossless beam splitters and phase rotations.
According to the Euler decomposition, any Gaussian unitary can be expressed as $\smash{U_{\rm p}\otimes_{j=1}^{N}U_{\text{sq},\lambda_{j}}\tilde{U}_{\rm p}}$, where $U_{\rm p}$ and $\tilde{U}_{\rm p}$ are passive multimode unitaries, and $U_{\text{sq},\lambda_{j}}$ are single-mode squeezing unitaries, described by $U_{\text{sq},\lambda_{j}}^{\dagger}(q_{j},p_{j})U_{\text{sq},\lambda_{j}}=(q_{j}e^{-\lambda_{j}},p_{j}e^{\lambda_{j}})$. Williamson’s theorem states that a Gaussian unitary can transform any Gaussian state into a thermal state. This implies that covariance matrices can be diagonalized using symplectic transformations, $\bm{\sigma}=\bm{S}\,\text{diag}(\nu_{1},\dots,\nu_{N},\nu_{1},\dots,\nu_{N})\bm{S}^{\top}$, where $\nu_{k}$ are known as the symplectic eigenvalues. Considering this theorem and noting that passive unitaries do not change the vacuum state, we can see that any pure Gaussian state can be constructed by applying single-mode squeezing unitaries on the vacuum state followed by a multimode passive Gaussian unitary. Note also that using the theorem one can see that the von Neumann entropy of Gaussian states can be expressed in terms of the symplectic eigenvalues $\sum_{k}\big{[}\frac{\nu_{k}+1}{2}\ln\!\left(\frac{\nu_{k}+1}{2}\right)-\frac{\nu_{k}-1}{2}\ln\!\left(\frac{\nu_{k}-1}{2}\right)\!\big{]}$.
Gaussian states can also be viewed as a thermal state of a quadratic Hamiltonian $\rho=e^{-\beta H}/{\rm tr}(e^{-\beta H})$ [38]. In this view, Williamson’s theorem implies that the Hamiltonian can be uncoupled by a Gaussian unitary $U^{\dagger}HU=\sum_{j=1}^{N}\omega_{j}(q_{j}^{2}+p_{j}^{2})$, where $\omega_{j}$ are the frequencies of the uncoupled modes, known as normal modes. Here, the Gaussian unitary $U$ corresponds to the symplectic transformation that diagonalizes the Hamiltonian matrix $\bm{S}^{\top}\bm{M}\bm{S}=\text{diag}(\omega_{1},\dots\omega_{N},\omega_{1},\dots\omega_{N})$. Using this Gaussian unitary, we can obtain the Hamiltonian’s ground state from the uncoupled Hamiltonian’s vacuum state.
Local Hamiltonians.
Let us first consider $N$ harmonic oscillators on a one-dimensional chain with the nearest-neighbor couplings and periodic boundary conditions. The general form of the Hamiltonian reads
$H_{\mu,\kappa}=\frac{1}{2}\sum_{n=1}^{N}\left[\frac{p_{n}^{2}}{\mu}+\mu m^{2}q_{n}^{2}+\kappa\left(q_{n+1}-q_{n}\right)^{2}\right].$ (2)
By using local squeezing operations $U_{\text{sq},\lambda}$ on each harmonic oscillator with $e^{2\lambda}=\sqrt{\mu\kappa}$ and defining $\epsilon=\sqrt{\mu/\kappa}$, this Hamiltonian can be transformed into
$H_{\epsilon}=\frac{1}{2}\sum_{n=1}^{N}\left[\frac{p_{n}^{2}}{\epsilon}+\epsilon\,m^{2}q_{n}^{2}+\frac{1}{\epsilon}\left(q_{n+1}-q_{n}\right)^{2}\right].$ (3)
This Hamiltonian known as the harmonic lattice model describes a Klein-Gordon (KG) free massive scalar field in two-dimensional spacetime, regularized on a lattice with the spacing $\epsilon$ and mass $m$. To recover the continuum limit, we have to take the limit of $\epsilon\to 0$. We, however, consider a non-vanishing lattice spacing, corresponding to a UV cut-off in the field theory description. For convenience, hereafter, we set $\epsilon=1$ and $H_{\mathrm{KG}}=H_{1}$.
The Hamiltonian matrix in (3) includes an $N\times N$ circulant matrix $\text{circ}(m^{2}+2,-1,0,\dots,0,-1)$, whose eigenvalues are $\smash{m^{2}+\big{(}2\sin({\pi(j-1)}/{N})\big{)}^{2}}$ for $j=1,\dots,N$. By ordering the eigenvalues $\smash{\omega_{1}^{2}\,{\leq}\,\omega_{2}^{2}\,{\leq}\cdots{\leq}\,\omega_{N}^{2}}$, an orthogonal matrix $\bm{V}$ that diagonalizes this matrix can be found as follows. Defining row vectors $\smash{F_{j}=\frac{1}{\sqrt{N}}(1,\nu^{(N+1-j)},\dots,\nu^{(N-1)(N+1-j)})}$ with $\nu=\exp(2\pi i/N)$, the first row of $\bm{V}$ is given by $V_{1}=F_{1}$, other rows are given by $V_{2k}=({F_{k+1}+F_{N-k+1}})/\sqrt{2}$ and $V_{2k+1}=i({F_{k+1}-F_{N-k+1}})/{\sqrt{2}}$ for $21\leq k<N/$ 2, and if $N$ is even the last row becomes $V_{N}=F_{\frac{N}{2}+1}$. Using this, the Hamiltonian matrix in (3) can be diagonalized by the symplectic matrix $\bm{S}_{\rm{p}}=\bm{V}\oplus\bm{V}$, corresponding to a passive Gaussian unitary $U_{\rm p}$. Applying this unitary on the Hamiltonian gives $U_{\rm p}^{\dagger}HU_{\rm p}=\sum_{j=1}^{N}(\omega_{j}^{2}q_{j}^{2}+p_{j}^{2})$. Therefore, by using additional local squeezing transformations $\otimes_{j=1}^{N}U_{\text{sq},\lambda_{j}}$ with squeezing parameters $\lambda_{j}=\frac{1}{2}\log\omega_{j}$, we obtain the Hamiltonian of uncoupled harmonic oscillators $\sum_{j=1}^{N}\omega_{j}^{2}(q_{j}^{2}+p_{j}^{2})$.
The ground state of Hamiltonian (3) is a Gaussian state of $N$ mode, which can be prepared in quantum optic settings by applying local squeezing operations $\otimes_{j=1}^{N}U_{\text{sq},\lambda_{j}}$ on the vacuum state of each mode, followed by an $N$-mode passive linear optical transformation described by orthogonal matrix $\bm{V}$. This analogy provides an operational cost of reaching the conformal field theory regime. Note that to reach this scale-invariant regime, we need to take the limit of $m\to 0$. However, since $\omega_{1}=m$, taking this limit implies infinite squeezing $\lambda_{1}\to-\infty$, meaning that preparing the ground state in the scale-invariant regime requires infinite energy.
Random Gaussian Unitary Dynamics.
We now describe our models of random Gaussian dynamics to investigate information scrambling. The first model (I) is a disordered version of the local Hamiltonian (3)
$H_{\mathrm{DKG}}=\frac{1}{2}\sum_{n=1}^{N}\left[\frac{p_{n}^{2}}{\epsilon}+\epsilon\,m^{2}q_{n}^{2}+\frac{J_{n}}{\epsilon}\left(q_{n+1}-q_{n}\right)^{2}\right].$ (4)
where $J_{n}$ are random real numbers. We denote this local Hamiltonian with random interactions by the disordered Klein-Gordon (DKG) Hamiltonian.
We also consider models in terms of passive linear-optical circuits to generate random Gaussian dynamics. The initial state is the tensor product of $N$ single-mode squeezed vacuum states, $\otimes_{j=1}^{N}U_{\text{sq},\lambda_{j}}|0,\dots,0\rangle$. In the first circuit model (IIa), a passive $N$-mode linear-optical network acts on the initial state in one step. The unitary operator $U_{\text{p}}$ linearly transforms the creation operators, $U_{\text{p}}^{\dagger}a^{\dagger}_{j}U_{\text{p}}=\sum_{n}\bm{U}_{jn}a^{\dagger}_{n}$, where $\bm{U}_{jn}$ are the elements of an $N\times N$ unitary transfer matrix $\bm{U}$, which describes the network. We choose matrix $\bm{U}$ from a Haar measure to induce randomness, and the corresponding symplectic matrix is given by $\bm{S}_{\text{p}}=\begin{pmatrix}\text{Re}(\bm{U})&\text{Im}(\bm{U})\\ -\text{Im}(\bm{U})&\text{Re}(\bm{U})\end{pmatrix}$. In the second circuit model (IIb), two layers of beam splitters described by $\otimes_{j=1}U_{\text{BS},2j}\otimes_{j=1}U_{\text{BS},2j-1}$ act on the $N$-mode state in each time step. We assume periodic boundary conditions, and the beam splitter operator $U_{\text{BS},j}$, which is the $2$-mode version of passive linear-optical networks, acts on $j$th and $(j+1)$th modes. We also choose the beam-splitter transfer matrix $\bm{U}_{\text{BS}}$ from a Haar measure to generate random dynamics.
In addition to the above random models, we consider the completely random model (IIIa), where the Hamiltonian matrix $\bm{M}$ in Eq. (1) is chosen from the Gaussian symplectic ensemble (GSE). We also consider another model (IIIb) with block diagonal Hamiltonian matrix $\bm{M}=\text{diag}((\bm{M}_{q}+\bm{M}_{q}^{\dagger})/2,\mathbb{1}_{N})$, where $\mathbb{1}_{N}$ is the $N\times N$ identity matrix and $\bm{M}_{q}$ is chosen from the Gaussian orthogonal ensemble (GOE) or the Gaussian unitary ensemble (GUE).
Figure 1: The disappearance of the memory effect, the dip in the time evolution of the joint entropy $S_{A_{1}\cup A_{2}}$, as an information scrambling diagnostic in the presence of randomness in the dynamics. Upper panel: $S_{A_{1}\cup A_{2}}$ in KG versus DKG models. The ground state of $H_{\mathrm{KG}}$ with $m=2$, denoted by $|\psi_{\text{KG},0}\rangle$ is evolved by unitary operators associated with $H_{\mathrm{KG}}$ with $m=10^{-7}$ (such a process is sometimes called a mass quench) and $H_{\mathrm{DKG}}$ with same value of $m$. We set the number of modes in the disjoint intervals $N_{A_{1}}=N_{A_{2}}=20$, the separation between them $d=60$, and the system size $N=500$. Lower panel: $S_{A_{1}\cup A_{2}}$ in the passive circuit model (IIb). The initial state is the tensor product of single-mode squeezed vacuum states with $\lambda_{i}=5$. We set $N_{A_{1}}=N_{A_{2}}=40$ and $d=200$. We consider two scenarios: the beam splitter (BS) transfer matrix $\bm{U}_{\mathrm{BS}}$ is chosen randomly at every single time step, and $\bm{U}_{\mathrm{BS}}$ is fixed during the evolution. For a balanced BS, we observe memory effects, where the second and third dips in the orange curve correspond to entanglement revivals. All random results are averaged over 500 samples.
Memory Effect During Entanglement Dynamics.
[…]
We consider the ground state of $H_{\mathrm{KG}}$ with $m\neq 0$, denoted by $|\psi_{\text{KG},0}\rangle$, as the initial state. The upper panel of Fig. 1 presents our simulation results for the joint entropy when the unitary evolution of $|\psi_{\text{KG},0}\rangle$ is induced by $H_{\mathrm{KG}}$ with $m\to 0$ (corresponding to free bosonic CFT with unit central charge), showing an apparent memory effect, versus $H_{\mathrm{DKG}}$ with the same $m$, showing no memory effect. We find similar behavior with the passive circuit model (IIb) presented in the lower panel. Note that in an integrable periodic system, the correlation between $A_{1}$ and $A_{2}$ revives periodically, known as entanglement revivals [46, 47, 48]. This effect can be observed in the lower panel of Fig. 1 for a balanced beam splitter with $\bm{U}_{\text{BS}}=(\mathbb{1}_{2}+i\bm{X})/\sqrt{2}$, where $\bm{X}$ is the $X$-Pauli matrix, while the memory effect disappears for the case of random beam splitters as another sign of scrambling in these systems.
Tripartite Mutual Information.
Tripartite mutual information is defined in terms of mutual information $I_{2}(A_{1}:A_{2})=S_{A_{1}}+S_{A_{2}}-S_{A_{1}\cup A_{2}}$ as
$I_{3}=I_{2}(A_{1}:A_{2})+I_{2}(A_{2}:A_{3})-I_{2}(A_{2}:A_{1}\cup A_{3}).$
The sign of $I_{3}$ is not definite in general [49, 50, 51] though in holographic theories $I_{3}\leq 0$ [25]. Negative TMI is considered a sign of information scrambling in dynamical systems [1, 2, 24]. Such a behavior has been realized in a range of many body systems [52, 53, 54, 55, 56].
Figure 2: Negative TMI in the presence of randomness in the dynamics as an indication of information scrambling. Upper panel: $I_{3}$ in KG versus DKG through a quenching process similar to Fig. 1. DKG results are averaged over 100 samples. We consider adjacent intervals $N_{A_{1}}=N_{A_{2}}=N_{A_{3}}=50$ and $N=1000$. Lower panel: $I_{3}$ for a uniform squeezed state ($\lambda_{i}=5$) evolved by a one-step $N$-mode Haar random unitary. We set $N=500$ and $\xi_{ij}\equiv\frac{N_{A_{i}}}{N_{A_{j}}}$.
We show that the sign of TMI for Gaussian states evolved by a unitary with some randomness takes a non-positive value. In the upper panel of Fig. 2, we show this for the initial state $|\psi_{\text{KG},0}\rangle$ evolved by unitary operators associated with $H_{\mathrm{DKG}}$ with $m=10^{-7}$. In the lower panel, we show the same effect for generic subregion configurations in our random model (IIb), where the tensor product of single-mode squeezed vacuum states is evolved by a Haar random passive linear-optical network. We have verified such a behavior in various other setups with different amounts of randomness.
Spectral Form Factor.
Compared to the nearest neighbor level statistics, SFF captures many particle level statistics. In our case in terms of the normal modes, SFF is defined for the $k$th mode in terms of the non-normalized single-mode partition function $Z_{k}(\beta)=\sum_{n_{k}}e^{-\beta\omega_{k}n_{k}}$ as
$g_{k}(\beta,t)$ $=\frac{|Z_{k}(\beta+it)|^{2}}{|Z_{k}(\beta)|^{2}}=\frac{\cosh(\beta\omega_{k})-1}{\cosh(\beta\omega_{k})-\cos(\omega_{k}t)},$
and SFF for the whole system is given by $g(\beta,t)=\Pi_{k}g_{k}(\beta,t)$. We numerically average $g(\beta,t)$ over thermal states, sometimes called as the quenched quantity.
Figure 3: We observe a dip followed by a non-linear ramp in SFF. We use $\beta^{-1}=500$ for the random model (IIIa) with $\bm{M}\in$ GSE, and $\beta^{-1}=1000$ for the random model (IIIb) with $\bm{M}_{q}\in$ GOE, GUE. The inset corresponds to the DKG Hamiltonian with $J_{n}\in(0.5,1.5)$. The system size is $N=500$ averaged over 100 samples.
While ergodic chaotic systems are known to exhibit a linear ramp at intermediate times, nonergodic integrable systems are not expected to show a ramp. Figure 3 shows our numerical results where we find a dip followed by a non-linear ramp for random models (IIIa) with $\bm{M}\in$ GSE and (IIIb) with $\bm{M}_{q}\in$ GOE, GUE. In all the cases, we see that the elevation of the non-linear ramp is proportional to the amount of randomness in the model. Note that an exponential ramp has been reported for SYK2 in [30, 31].
Out of Time-Ordered Correlators.
OTOC has been defined by quantizing the classical observation that the sensitivity to the initial conditions can be quantified by the Poisson bracket $\{q(t),p\}$ [21, 3, 57]. Although the old paradigm expresses the exponential growth of OTOC and quantum chaos as reciprocal notions [3], in the current understanding an unstable saddle point in integrable systems can lead to the exponential growth of OTOC [27]. A natural question is how OTOC behaves in integrable systems in the presence of randomness.
Here we consider OTOC in terms of the canonical operators
$C_{\beta,jk}(t)=-{\rm tr}\!\left(\frac{e^{-\beta H}}{{\rm tr}\left(e^{-\beta H}\right)}\big{[}q_{j}(t),p_{k}\big{]}^{2}\right)$ (5)
where $q_{j}(t)=e^{itH}q_{j}e^{-itH}$ and $j,k$ stands for the mode index. For the cases of KG, DKG, our random model (IIIb) as well as any Hamiltonian matrix that can be diagonalized with a symplectic transformation of the form $\bm{S}=\bm{S}_{q}\oplus\bm{S}_{p}$, we find (see Appendix A)
$\begin{split}C_{\infty,jk}(t)&=\big{(}\bm{S}_{q}\,\mathrm{diag}(\cos(\omega_{1}t),\cdots,\cos(\omega_{N}t))\bm{S}_{p}^{\top}\big{)}_{jk}^{2},\end{split}$ (6)
where $\beta\to\infty$ implies zero temperature and the OTOC is with respect to the ground state of the Hamiltonian. In this case, using $\bm{S}_{q}\bm{S}_{p}^{\top}=\mathbb{1}_{N}$, one can easily verify that $C_{\infty,jk}(t\ll 1)\sim t^{4}$ for $k\neq j$ and $C_{\infty,jk}(t\ll 1)\sim t^{2}$ for $k=j$. Our detailed analytic analysis and the numerical simulation shown in Fig. 4 perfectly agree with this result.
In Appendix B, we have analyzed a generic form of OTOC in bosonic systems showing that it merely depends on a single time-dependent parameter, which is a function of the normal modes frequencies and the random unitary. Furthermore, we interestingly show that the saturation time decreases with the number of oscillators as $t_{*}\sim 1/\log N$. This feature results in shortening the power law growth in large systems followed by a plateau.
Figure 4: OTOC for $q_{j}(t)$ and $p_{k}$ in the ground state of the KG, DKG and model (IIIb) Hamiltonians, showing a power law growth. We set $j=1$, $k=N/2$, and $N=100$. The DKG and GOE results are averaged over 100 samples. The dashed red line $\propto t^{4}$.
Discussions.
We have shown that randomness in the couplings of integrable bosonic systems features scrambling measures to behave as those in chaotic systems. Randomness in such systems destroys the memory effect in entanglement spread and results in negative TMI similar to holographic states. Furthermore, it causes the SFF to show a (non-linear) ramp but OTOC in these models cannot grow faster than a power law, in which the duration is shortened by increasing the number of modes.
Our study should not be mixed with what has been addressed as weak scrambling in nonrandom local integrable systems [58, 20]. There is no global scrambling effect in such systems but rather a local effect in the sense that recovering the initial data of a subregion would need measurements in a larger subregion. Such an effect is enhanced in Lifshitz models due to non-linear dispersion which results in infinitely prolonged saturation time for the entanglement of finite subregions [59].
Our results suggest possible semiclassical descriptions for states in random integrable systems which deserve further investigation. Among several open questions about random integrable models, an interesting one would be the description of such dynamical behaviors in terms of ballistically propagating quasiparticles.
Acknowledgments:
We thank Reza Mohammadi Mozaffar and Pratik Nandy for useful discussions. AM would like to acknowledge support from ICTP through the Associates Programme (2023-2028) and for hospitality during stages of this work.
Appendix A Appendix A: Derivation of out-of-time-order correlators for quadrature operators
We derive Eq. (6), OTOC for canonical operators, and the ground state of quadratic Hamiltonians whose matrix $\bm{M}$ can be diagonalized using symplectic transformations of the form $\bm{S}=\bm{S}_{q}\oplus\bm{S}_{p}$. This includes KG, DKG, and our random model (IIIb) Hamiltonians. We define
$H_{\rm D}=U^{\dagger}HU=\frac{1}{2}U^{\dagger}r^{\top}U\bm{M}U^{\dagger}rU=\frac{1}{2}r^{\top}\bm{D}r$ (7)
where we used $U^{\dagger}rU=\bm{S}r$ and matrix $\bm{D}=\bm{S}^{\top}\bm{M}\bm{S}$ is diagonal.
Assuming that $\bm{S}=\bm{S}_{q}\oplus\bm{S}_{p}$, we have
$\Big{[}e^{iH_{\rm D}t}U^{\dagger}q_{j}Ue^{-iH_{\rm D}t},U^{\dagger}p_{k}U\Big{]}$ $=\sum_{n,m}\bm{S}_{q,jn}\bm{S}_{p,km}\Big{[}e^{iH_{\rm D}t}q_{n}e^{-iH_{\rm D}t},p_{m}\Big{]}$
$=\sum_{n,m}\bm{S}_{q,jn}\bm{S}_{p,km}\big{[}q_{n}\cos(\omega_{k}t)+p_{n}\sin(\omega_{k}t),p_{m}\big{]}$
$=i\sum_{n}\bm{S}_{q,jn}\bm{S}_{p,kn}\cos(\omega_{n}t)$
$=i\Big{(}\bm{S}_{q}\,\text{diag}\big{(}\cos(\omega_{1}t),\cdots,\cos(\omega_{N}t)\big{)}\bm{S}_{p}^{\top}\Big{)}_{ij},$
where we used $e^{iH_{\rm D}t}q_{n}e^{-iH_{\rm D}t}=q_{n}\cos(\omega_{k}t)+p_{n}\sin(\omega_{k}t)$ in the second line, and $[q_{n},p_{m}]=i\delta_{nm}$ in the third line. Using the above relation and $H=UH_{\rm D}U^{\dagger}$, OTOC for canonical operators reads
$C_{\beta,jk}(t)$ $=-{\rm tr}\!\left(\frac{e^{-\beta H}}{{\rm tr}\left(e^{-\beta H}\right)}\big{[}q_{j}(t),p_{k}\big{]}^{2}\right)$
$=-{\rm tr}\left(\frac{e^{-\beta H_{\rm D}}}{{\rm tr}\left(e^{-\beta H_{\rm D}}\right)}\Big{[}e^{iH_{\rm D}t}U^{\dagger}q_{j}Ue^{-iH_{\rm D}t},U^{\dagger}p_{j}U\Big{]}^{2}\right)$
$=\Big{(}\bm{S}_{q}\,\text{diag}\big{(}\cos(\omega_{1}t),\cdots,\cos(\omega_{N}t)\big{)}\bm{S}_{p}^{\top}\Big{)}_{ij}^{2},$ (8)
which is Eq. (6) in the main text. This expression can be viewed as a generalization of the same quantity for a single oscillator previously reported in [60].
Appendix B Appendix B: Out-of-time-order correlators for generic bosonic operators
Considering an $N$-mode bosonic system with the Hamiltonian $H$ and local operators $W_{j}$ and $V_{k}$ acting on the $j$th and $k$th modes, respectively, OTOC can alternatively be defined as
$F_{\beta,jk}(t)=\left\langle W^{\dagger}_{j}(t)V^{\dagger}_{k}W_{j}(t)V_{k}\right\rangle={\rm tr}\Bigg{(}\!\frac{e^{-\beta H}}{{\rm tr}(e^{-\beta H})}W^{\dagger}_{j}(t)V^{\dagger}_{k}W_{j}(t)V_{k}\!\Bigg{)},$ (9)
where $W_{j}(t)=e^{-itH}W_{j}e^{itH}$. Note that if $W_{j}$ and $V_{k}$ are unitary, then the relation between this quantity and $C_{\beta,jk}(t)=\big{\langle}[V_{k},W_{j}(t)]^{\dagger}[V_{k},W_{j}(t)]\big{\rangle}$, which is used in the main text, is given by $\text{Re}\big{[}F_{\beta,jk}(t)\big{]}=1-C_{\beta,jk}(t)/2$. For simplicity, we consider Hamiltonians that can be transformed into an uncoupled Hamiltonian of harmonic oscillators using a passive Gaussian unitary,
$H=(a_{1},\dots,a_{N})\bm{M}(a_{1}^{\dagger},\dots,a_{N}^{\dagger})^{\top}=U_{\text{p}}H_{\text{D}}U_{\text{p}}^{\dagger}$ (10)
with $H_{\text{D}}=\sum_{j}\omega_{j}a^{\dagger}_{j}a_{j}$. The passive unitary can be described by $U_{\text{p}}^{\dagger}a^{\dagger}_{j}U_{\text{p}}=\sum_{n}\bm{U}_{jn}a^{\dagger}_{n}$ with $\bm{U}_{jn}$ being the elements of the $N\times N$ unitary matrix that diagonalizes the Hamiltonian matrix, $\sum_{jk}\bm{U}^{*}_{mk}\bm{M}_{kj}\bm{U}_{jn}=\omega_{n}\delta_{nm}$. By using this transformation, we have
$F_{\beta,jk}(t)={\rm tr}\Bigg{(}\!\frac{e^{-\beta H_{\text{D}}}}{{\rm tr}(e^{-\beta H_{\text{D}}})}\tilde{W}^{\dagger}_{j}(t)\tilde{V}^{\dagger}_{k}\tilde{W}_{j}(t)\tilde{V}_{k}\!\Bigg{)}$ (11)
where $\tilde{W}_{j}(t)=e^{itH_{\text{D}}}U_{\text{p}}^{\dagger}W_{j}U_{\text{p}}e^{-itH_{\text{D}}}$ and $\tilde{V}_{k}=U_{\text{p}}^{\dagger}V_{k}U_{\text{p}}$.
The thermal state of $N$-mode uncoupled system can be expressed in terms of multimode coherent states $|\alpha_{1},\dots,\alpha_{N}\rangle$ using the Glauber-Sudarshan representation [61, 62]
$\frac{e^{-\beta H_{\text{D}}}}{{\rm tr}(e^{-\beta H_{\text{D}}})}=\idotsint d^{2}\alpha_{1}\dots d^{2}\alpha_{N}\frac{e^{-|\alpha_{1}|^{2}/\bar{n}_{1}}}{\bar{n}_{1}\pi}\times\dots\times\frac{e^{-|\alpha_{N}|^{2}/\bar{n}_{N}}}{\bar{n}_{N}\pi}|\alpha_{1},\dots,\alpha_{N}\rangle\langle\alpha_{1},\dots,\alpha_{N}|$ (12)
with $\bar{n}_{j}=(e^{\beta\omega_{j}}-1)^{-1}$ being the mean photon number in the $j$th mode. Using this, we can write
$F_{\beta,jk}(t)=\int d^{2N}\!\bm{\alpha}\;P(\bm{\alpha})\tilde{F}_{\bm{\alpha},jk}(t),$ (13)
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{N})$, $P(\bm{\alpha})=\prod_{j=1}^{N}\exp\!\big{(}\!-|\alpha_{j}|^{2}/\bar{n}_{j}\big{)}/(\bar{n}_{j}\pi)$, and
$\tilde{F}_{\bm{\alpha},jk}(t)=\big{\langle}\alpha_{1},\dots,\alpha_{N}\big{|}\tilde{W}^{\dagger}_{j}(t)\tilde{V}^{\dagger}_{k}\tilde{W}_{j}(t)\tilde{V}_{k}\big{|}\alpha_{1},\dots,\alpha_{N}\big{\rangle}$ (14)
is the OTOC in terms of coherent states. Note that for $\beta=\infty$ (zero temperature) $F_{\infty,jk}(t)=\tilde{F}_{\bm{0},jk}(t)$.
The operators can also be expanded in terms of displacement operators $D_{j}(\zeta)=\exp(\zeta a^{\dagger}_{j}-\zeta^{*}a_{j})$ with $\zeta\in\mathbb{C}$ [63]
$W_{j}=\frac{1}{\pi}\int d^{2}\zeta\;\Phi_{W_{j}}\!(\zeta)\;D_{j}(-\zeta),$ (15)
where $\Phi_{W_{j}}\!(\zeta)={\rm tr}\big{(}W_{j}D_{j}(\zeta)\big{)}$ is the characteristic function of operator $W_{j}$. We have
$U_{\text{p}}^{\dagger}D_{j}(\zeta)U_{\text{p}}=D_{1}(\zeta\bm{U}_{j1})\otimes\dots\otimes D_{N}(\zeta\bm{U}_{jN}),$ (16)
where we used $U_{\text{p}}^{\dagger}a^{\dagger}_{j}U_{\text{p}}=\sum_{n}\bm{U}_{jn}a^{\dagger}_{n}$, and also
$e^{itH_{\text{D}}}D_{j}(\zeta)e^{-itH_{\text{D}}}=D_{j}(\zeta e^{it\omega_{j}}).$ (17)
Using these two relations we can then compute
$\tilde{W}_{j}(t)=e^{itH_{\text{D}}}U_{\text{p}}^{\dagger}W_{j}U_{\text{p}}e^{-itH_{\text{D}}}=\frac{1}{\pi}\int d^{2}\zeta\;\Phi_{W_{j}}\!(\zeta)D_{1}(-\zeta\bm{U}_{j1}e^{it\omega_{1}})\otimes\dots\otimes D_{N}(-\zeta\bm{U}_{jN}e^{it\omega_{N}}).$ (18)
Therefore, by using a similar expression for $\tilde{V}_{k}$, the OTOC in terms of coherent states can be written as
$\begin{split}\tilde{F}_{\bm{\alpha},jk}(t)=\frac{1}{\pi^{4}}&\int d^{2}\zeta d^{2}\xi d^{2}\beta d^{2}\gamma\;\Phi^{*}_{W_{j}}\!(\zeta)\Phi^{*}_{V_{k}}\!(\xi)\Phi_{W_{j}}\!(\beta)\Phi_{V_{k}}\!(\gamma)\\ &\times\prod_{n=1}^{N}\big{\langle}\alpha_{n}\big{|}D_{n}(\zeta\bm{U}_{jn}e^{it\omega_{n}})D_{n}(\xi\bm{U}_{kn})D_{n}(-\beta\bm{U}_{jn}e^{it\omega_{n}})D_{n}(-\gamma\bm{U}_{kn})\big{|}\alpha_{n}\big{\rangle}.\end{split}$ (19)
As we can see each term in the product oscillates with time with frequency $\omega_{n}$. By using the following relation for displacement operators
$D_{n}(\mu)D_{n}(\nu)=\exp\!\bigg{(}\!\frac{\mu\nu^{*}-\nu\mu^{*}}{2}\!\bigg{)}D_{n}(\mu+\nu),$ (20)
we get
$\begin{split}D_{n}(\zeta\bm{U}_{jn}&e^{it\omega_{n}})D_{n}(\xi\bm{U}_{kn})D_{n}(-\beta\bm{U}_{jn}e^{it\omega_{n}})D_{n}(-\gamma\bm{U}_{kn})=D_{n}\big{(}\bm{U}_{jn}e^{it\omega_{n}}(\zeta-\beta)+\bm{U}_{kn}(\xi-\gamma)\big{)}\\ &\times\exp\bigg{[}\frac{1}{2}\Big{(}(\zeta\xi^{*}+\beta\gamma^{*}+\beta\xi^{*}-\zeta\gamma^{*})\bm{U}_{jn}e^{i\omega_{n}t}\bm{U}_{kn}^{*}+\gamma\xi^{*}|\bm{U}_{kn}|^{2}+\beta\zeta^{*}|\bm{U}_{jn}|^{2}\Big{)}-\text{c.c.}\bigg{]},\end{split}$ (21)
where c.c. stands for complex conjugate. We also have $\langle\alpha|D(\mu)|\alpha\rangle=\exp(-|\mu|^{2}/2+\mu\alpha^{*}-\alpha\mu^{*})$. Using these equations, Eq. (19) becomes
$\begin{split}\tilde{F}_{\bm{\alpha},jk}(t)=\frac{1}{\pi^{4}}&\int d^{2}\zeta d^{2}\xi d^{2}\beta d^{2}\gamma\;\Phi^{*}_{W_{j}}\!(\zeta)\Phi^{*}_{V_{k}}\!(\xi)\Phi_{W_{j}}\!(\beta)\Phi_{V_{k}}\!(\gamma)\\ &\times\exp\!\bigg{[}\!-\frac{1}{2}|\zeta-\beta|^{2}-\frac{1}{2}|\xi-\gamma|^{2}-\frac{1}{2}\big{(}(\zeta-\beta)(\xi^{*}-\gamma^{*})\bm{V}_{jk}(t)+\text{c.c.}\big{)}\bigg{]}\\ &\times\exp\bigg{[}\Big{(}(\zeta-\beta)\sum_{n=1}^{N}\bm{U}_{jn}e^{it\omega_{n}}\alpha_{n}^{*}+(\xi-\gamma)\sum_{n=1}^{N}\bm{U}_{kn}\alpha_{n}^{*}\Big{)}-\text{c.c.}\bigg{]}\\ &\times\exp\!\bigg{[}\frac{1}{2}\Big{(}(\zeta\xi^{*}+\beta\gamma^{*}+\beta\xi^{*}-\zeta\gamma^{*})\bm{V}_{jk}(t)+\gamma\xi^{*}+\beta\zeta^{*}\Big{)}-\text{c.c.}\bigg{]},\end{split}$ (22)
where we used $\sum_{n}|\bm{U}_{jn}|^{2}=1$ and
$\bm{V}_{jk}(t)=\sum_{n=1}^{N}\bm{U}_{jn}e^{i\omega_{n}t}\bm{U}_{kn}^{*}$ (23)
is an element of the unitary matrix $\bm{V}(t)=\exp(i\bm{M}t)$, where $\bm{M}$ is the Hamiltonian matrix in Eq. (10). Given the characteristic functions $\Phi_{V_{k}}$ and $\Phi_{W_{j}}$, one can therefore compute $F_{\bm{\alpha},jk}(t)$ and then by averaging with respect to $P(\bm{\alpha})$ in Eq. (13), $F_{\beta,jk}(t)$ can be obtained.
As a simple example, suppose that $W_{j}=D_{j}(\mu)$ and $V_{k}=D_{k}(\nu)$ are displacement operators with the characteristic functions $\Phi_{W_{j}}(\beta)=\pi\delta^{2}(\beta+\mu)$ and $\Phi_{V_{k}}(\gamma)=\pi\delta^{2}(\gamma+\nu)$. In this case, we have
$\tilde{F}_{\bm{\alpha},jk}(t)=\exp\!\big{(}\mu\nu^{*}\bm{V}_{jk}(t)-\mu^{*}\nu\bm{V}^{*}_{jk}(t)\big{)},$ (24)
which implies that $F_{\beta,jk}(t)=\tilde{F}_{\bm{\alpha},jk}(t)$ is independent of $\beta$. This also gives $C_{\beta,jk}(t)=2-2\text{Re}[F_{\beta,jk}(t)]=4\sin^{2}\!\big{(}\text{Im}[\mu\nu^{*}\bm{V}_{jk}(t)]\big{)}$.
In this example, we can see that the time-dependent behavior of OTOC is determined by that of $\bm{V}_{jk}(t)$. Therefore, to further investigate the time-dependent behaviour of $\bm{V}_{jk}(t)$, we ran numerical simulations for a randomly selected set of $\{\omega_{n}\}$ and $\bm{U}$ selected as a $N\times N$ unitary matrix chosen from the Haar measure. As shown in the left panel of Fig. 5, $\bm{V}_{jk}(t)$ rapidly saturates to a constant value at some time $t_{*}\sim 1/\log N$. The larger the system size, the faster $\bm{V}_{jk}(t)$ and the OTOC saturation. Note that this behavior is critically different from the scrambling time $t_{s}\sim\log N$ in generic chaotic systems. Due to this rapid saturation, OTOC does not have enough time to oscillate except for the cases where the parameters in $W_{j}=D_{j}(\mu)$ and $V_{k}=D_{k}(\nu)$ slow down this rapid saturation time (see the inset of the right panel in Fig. 5).
Figure 5: Right: Time evolution of the diagonal and off-diagonal elements of $\bm{V}_{jk}(t)$. $\bm{U}$ is a Haar random unitary matrix and the result is averaged over 200 samples. Left: The behavior of OTOC $F_{\beta,jk}(t)$ for displacement operators with parameters $\mu$ and $\nu$. The inset shows the choice of $\mu\nu^{*}\sim 1/N^{2}$ is small enough to compensate for the rapid saturation of $\bm{V}_{jk}(t)$ and allow the OTOC to oscillate in time.
Another example is to consider canonical operators as local operators $W_{j}=q_{j}=(a_{j}+a_{j}^{\dagger})/\sqrt{2}$ and $V_{k}=p_{k}=-i(a_{k}-a_{k}^{\dagger})/\sqrt{2}$ whose characteristic functions are $\Phi_{W_{j}}(\beta)=\pi(\frac{\partial}{\partial\beta}-\frac{\partial}{\partial\beta^{*}})\delta^{2}(\beta)/\sqrt{2}$ and $\Phi_{V_{k}}(\gamma)=i\pi(\frac{\partial}{\partial\gamma}+\frac{\partial}{\partial\gamma^{*}})\delta^{2}(\gamma)/\sqrt{2}$. Plugging these into (22) for $\bm{\alpha}=\bm{0}$ and using integration by parts yields
$\tilde{F}_{\mathbf{0},jk}(t)=F_{\infty,jk}(t)=\frac{1}{4}\left(1+|\bm{V}_{jk}(t)|^{2}-{\bm{V}^{*}_{jk}}^{2}(t)\right)$ (25)
which is again governed by the behaviour of $\bm{V}_{jk}(t)$. Using the above formalism, it is straightforward to show that the other definition of OTOC at zero temperature is given by
$C_{\infty,jk}(t)=\frac{1}{4}\left(2|\bm{V}_{jk}(t)|^{2}+\bm{V}_{jk}^{2}(t)+{\bm{V}^{*}_{jk}}^{2}(t)\right)=\big{[}\mathrm{Re}(\bm{V}_{jk}(t))\big{]}^{2}\,.$ (26)
As shown in Fig. 6, this function shows a power law growth, similar to the behavior in Fig. 5. Note that for a single oscillator, this is a periodic function (see Eq. (6) and [60]) while for $N>1$ the role of $\bm{V}_{jk}(t)$ becomes important.
Figure 6: The behavior of OTOC $C_{\infty,jk}(t)$ for the canonical operators $V=q_{j}$ and $W=p_{k}$. The results are averaged over 300 samples.
The above formalism can also be applied to compute OTOC for $q_{j}^{n_{q}}$ and $p_{k}^{n_{p}}$. In this case, the characteristic functions are given in terms of higher-order partial derivatives of the Dirac delta function (22) that results in an integer power of $\mathrm{Im}(\bm{V}_{jk}(t))$ and $\mathrm{Re}(\bm{V}_{jk}(t))$ in OTOC. Note that, in general, due to the rapid saturation of $\bm{V}_{jk}(t)$, the result of OTOC reduces to a power law in $t$.